OpenStudy (anonymous):

FAN AND MEDAL!!!!! Which of the following expressions is the inverse of the function y equals quantity x plus 5 divided by 7

OpenStudy (anonymous):

answer choices y= x+7 divide by 5 y=x-5 divide by 7 y= 5x-7 y=7x-5

OpenStudy (anonymous):

i think it is d

OpenStudy (ksaimouli):

f(x)=(x+5)/7

OpenStudy (ksaimouli):

solve for x, you get= 7y-5

OpenStudy (ksaimouli):

yes you are right

OpenStudy (ksaimouli):

it is d